A torch bulb is rated as 2.5V and 600 ma. Calculate its the energy consumed if the bulb is lighted for 4 hours?
1
Expert's answer
2013-09-25T12:48:10-0400
how much does it consumes in 1 second?

P = U*I = 2.5 V* 0.6 A = 1.5 J/s

How many seconds in 4 hours?
3600*4 = 14400 s

Hence, i 4 hours it will consume

1.5*14400 = 21600 J